OpenStudy (anonymous):

What is the simplified form of (3/x^2) / (1/x^3)

OpenStudy (anonymous):

dividing fractions is like multiple the inverse : (3/x^2) * (x^3/1)

OpenStudy (anonymous):

so can you solve it now ?

OpenStudy (anonymous):

3x^3/x^2

OpenStudy (anonymous):

but then how would you get the restrictions

OpenStudy (anonymous):

ok .. but you can go further .. using : x^m/x^n = x^(m-n)

OpenStudy (anonymous):

so your expression will become : 3x^(3-2) = 3x

OpenStudy (anonymous):

but.. you must remember that when you started there was x^3 in the denominator so x is not 0

OpenStudy (anonymous):

x^3 and x^2 in the denominators

OpenStudy (anonymous):

im so confused so would the simplified form be just 3x?

OpenStudy (anonymous):

friend see my attachment i hope it helps.

OpenStudy (anonymous):

so i was right?? It is 3x?!

OpenStudy (anonymous):

yes friend you were right.

OpenStudy (anonymous):

yaaay thank you:)